### v1

@@ 8.1 Classification Metrics
In multi-class HGR problems—especially when dealing with class imbalance—researchers often report both micro-averaged (per-instance) and macro-averaged (per-class) versions of common classification metrics. Micro-averaged metrics aggregate contributions from all instances across classes, thus being sensitive to class imbalance. Macro-averaged metrics compute the metric independently for each class and then average, treating all classes equally regardless of their frequency.
Micro-averaged metrics:
$\text{Accuracy}_{micro}=\frac{\sum_{i=1}^{N}(TP_{i}+TN_{i})}{\sum_{i=1}^{N}(TP_{i}+TN_{i}+FP_{i}+FN_{i})}$ (1)
$\text{Precision}_{micro}=\frac{\sum_{i=1}^{N}TP_{i}}{\sum_{i=1}^{N}(TP_{i}+FP_{i})}$ (2)
$\text{Recall}_{micro}=\frac{\sum_{i=1}^{N}TP_{i}}{\sum_{i=1}^{N}(TP_{i}+FN_{i})}$ (3)
[…]
Macro-averaged metrics:
$\text{Accuracy}_{macro}=\frac{1}{N}\sum_{i=1}^{N}\frac{TP_{i}+TN_{i}}{TP_{i}+FP_{i}+TN_{i}+FN_{i}}$ (5)
$\text{Precision}_{macro}=\frac{1}{N}\sum_{i=1}^{N}\frac{TP_{i}}{TP_{i}+FP_{i}}$ (6)
$\text{Recall}_{macro}=\frac{1}{N}\sum_{i=1}^{N}\frac{TP_{i}}{TP_{i}+FN_{i}}$ (7)
[…]
F1-Score: For both micro and macro cases, the F1-score is defined as the harmonic mean of Precision and Recall:
$F_{1}=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$ (9)
Here, $TP_{i}$, $FP_{i}$, $TN_{i}$, and $FN_{i}$ denote the number of true positives, false positives, true negatives, and false negatives for the $i$-th class, respectively, and $N$ is the total number of classes.
@@ 8.2 Sequence-to-Sequence Metrics
For continuous HGR (CHGR), particularly in Continuous Sign Language Recognition (CSLR), sequence-level metrics are used. Word Error Rate (WER) [122, 2] is the standard metric, measuring the minimum number of substitutions ($S$), deletions ($D$), and insertions ($I$) needed to transform the predicted sequence into the ground truth sequence of length $N$:
@@
Edit Distance Accuracy [45, 185] is a complementary metric representing the ratio of correctly predicted elements in the sequence:
[…]
@@ 8.3 Computational Resource Metrics
  • Model Size: The number of trainable parameters or total memory footprint of the model.
@@
8.4 Overfitting Indicators
To assess generalization and detect overfitting, the following indicators are used:
  • Train-Test Performance Gap: A significant difference between training and testing accuracy (or loss) often signals overfitting.
  • Validation Loss Trend: Increasing validation loss while training loss decreases is a common overfitting pattern.
  • Early Stopping Criteria: Monitored via a plateau or increase in validation loss across epochs.
@@ 9 Current Challenges and Future Research Directions
As mentioned in the Methods sections (Sections 4, 5, and 6), researchers have developed sophisticated techniques and large-scale datasets to achieve satisfactory performance. However, as shown in the relevant tables (Tables 3, 4, 5, 6, 7), recognition accuracy remains low in many cases—particularly in more complex tasks such as Continuous Sign Language Recognition (CSLR). In this section, we discuss the current challenges and limitations that hinder efficient Human Gesture Recognition (HGR), posing ongoing concerns for the research community and shaping future research directions. Additionally, we review the techniques proposed to address these issues, offering insights that may guide future efforts to overcome the identified challenges.
@@ 9.1 Gesture-Irrelevant Factors
The presence of gesture-irrelevant factors (or redundant information) hinders accurate recognition in diverse real-world scenarios. Some of these factors include the following [171, 9, 80, 17, 141, 206]:
@@
Aiming to address these challenges, researchers have proposed segmentation models to crop hand regions and remove background interference [208, 34, 153], object detectors for spatial normalization [17, 9], super-resolution techniques to enhance image quality [17], attention mechanisms [206, 207] to enforce the visual encoder to pay attention on the most informative regions, information-theoretic losses [100], and feature selection techniques [173] to reduce information redundancy. Many of these approaches rely on computationally expensive off-the-shelf models, which increase the complexity of the overall HGR methods and set an upper bound to their accuracy. Thus, the presence of gesture-irrelevant factors could concern future research work, while potentially being combined with eXplainable Artificial Intelligence (XAI) approaches [148] to facilitate this task.
@@ 9.2 Data Scarcity
Many existing datasets either contain a relatively small number of samples or lack sufficient diversity. For example, among static HGR datasets, only the versions of HaGRID [86, 131] have been collected from more than one hundred participants, despite some others having a plausible number of samples [67, 17]. A similar situation is observed in IDHGR datasets, where only the Jester dataset [117] includes data from over one hundred subjects. Furthermore, the amount of samples for specific modalities, such as IR data, is limited compared with the size of RGB-based datasets, an issue known as data imbalance [181]. The lack of large-scale, diverse datasets leads to overfitting and limited generalization, which is closely related to the challenges discussed earlier. This issue is further exacerbated in tasks that involve costly annotation processes and require data-intensive deep learning models with a large number of parameters, such as CSLR.
@@
  • Multilingual Data Aggregation: Combining datasets (e.g., CSL and DGS) to exploit cross-lingual visual similarities increases the overall amount of training data [188].
  • Self-Supervised Learning: In the context of the self-supervised learning setting [91], gloss-to-gloss translation tasks have been used for parameter initialization [205], providing linguistic priors to sign language models. Pretraining with pose reconstruction has also been investigated to incorporate prior knowledge of hand structure [71, 68, 79].
  • Advanced Data Augmentation: Adversarial learning has been employed to optimize augmentation hyperparameters [128].
  • Few-/Zero-shot Learning: Approaches using only a small number of samples to correctly recognize gestures—often based on CLIP—have also been proposed [216, 24, 147, 146, 22].
These approaches have led to improved recognition performance, even when trained on limited data. However, accuracy remains suboptimal, making data scarcity a persistent research challenge and a key direction for future work in the field.
@@ 9.3 Optimization Difficulties
Training the visual encoder effectively remains difficult, especially in deep architectures for CSLR. As shown in CTCA [57] (Fig. 17), increasing the depth of the temporal module (e.g., 1D-TCN followed by BiLSTM) improves its generalization but hampers the encoder’s learning due to gradient attenuation in backpropagation. Furthermore, the CTC spike phenomenon—where the model over-focuses on a few keyframes—can cause rapid overfitting and limit encoder training [122, 62].
Figure 17: Experimental results of a study conducted by authors in [57], which aims to measure the impact of the depth of the temporal aggregation module (a 1D-TCN with various number of layers followed by a two-layer BiLSTM) on the effectiveness of training the visual encoder (aka spatial perception module). In both diagrams, the x axis represents the number of layers of the 1D-TCN. In the top chart, the two modules are compared considering the information stored in weights (IIW) [187] (y axis), which can be viewed as an indicator of generalization gap, while in the bottom the corresponding WER is shown. As the number of layers of the 1D-TCN increases, although the generalization ability of the short-term temporal module increases, the overall performance decreases due to the insufficient training of the visual encoder.
@@
  • Auxiliary Supervision: Adding early-stage classifiers to the encoder or short-term module has been proposed as a strategy to guide training by providing additional supervision [122, 143, 74, 76, 111, 213].
  • Temporal Module Redesign: Placing the 1D-TCN and BiLSTM in parallel instead of sequentially [57].
  • Linguistic Constraints and Gloss Segmentation: Using gloss boundaries [62] or prior knowledge [58] to regularize learning.
The aforementioned approaches —particularly auxiliary supervision [122]— have served as the foundation for many existing methods. However, the increasing complexity of modern deep architectures often exceeds the effectiveness of these techniques. As a result, optimization remains a significant challenge —one that may be further exacerbated by the continued development of even deeper models.
@@ 9.4 Limitations of Architectures
Despite the achievements of DL, the existing architectures suffer from the following limitations:
  • Block-fixed models: The existing architectures are based on hand-crafted design, which may not be optimal[206].
  • GCNs with fixed topology graphs: Standard GCNs use fixed topology graphs based on hand anatomy, which may not capture optimal inter-joint relationships.
  • Ineffective multimodal fusion: Most of the current works utilize late fusion, that does not fully leverage the available information [68, 71, 202, 204, 203, 104, 82, 81].
To address these limitations, researchers have proposed various strategies, including Neural Architecture Search (NAS) to identify optimal model structures [206], learnable or adaptive skeleton adjacency matrices [77, 30, 28, 107], and multi-level fusion for dual-stream approaches [60]. Despite these advances, several aspects of deep learning architectures remain underexplored, leaving a substantial research gap.
@@ 10 Conclusions
The undeniable importance of vision-based hand gesture recognition (HGR) stems from its wide range of applications—spanning human-robot interaction to sign language understanding. This broad applicability has made it a vibrant and active research field. Recent advances in deep learning have played a pivotal role in driving progress, enabling the development of increasingly sophisticated architectures and training strategies aimed at tackling more complex and fine-grained tasks, such as continuous sign language recognition.
This survey focused on HGR methods that leverage visual input and deep learning. These methods were highlighted not only for their superior performance compared to traditional hand-crafted approaches, but also for their higher potential for deployment in real-world scenarios, unlike more cumbersome sensor-based techniques. From the surveyed literature, it becomes clear that even relatively simple and shallow deep learning models can achieve high accuracy in static HGR tasks. In such cases, the primary challenges are typical of general computer vision problems and are not inherently tied to HGR itself.
Despite the significant progress achieved in vision-based hand gesture recognition, several open challenges remain that shape future research directions. A major concern is the generalization of models across diverse environments, lighting conditions, camera viewpoints, and user-specific variations, which often lead to performance drops in real-world settings. The lack of large-scale, diverse, and well-annotated datasets, especially for continuous gestures and sign language, also hinders the development of more robust models. Furthermore, model efficiency and scalability are critical for deployment on resource-constrained edge devices, necessitating advances in lightweight architectures and model compression techniques. The integration of multimodal data, such as combining vision with audio or inertial signals, remains underexplored yet promising. Finally, interpretability and trustworthiness of deep learning models are essential, particularly for safety-critical applications like healthcare or assistive technologies. Addressing these challenges will not only enhance the practical applicability of HGR systems but also contribute to more inclusive and accessible human-computer interaction paradigms.
[…]
Acknowledgments The research leading to these results received funding from the European Commission under Grant Agreement No. 101168042 (TRIFFID) and No. 101189557 (TORNADO).

### v2

@@ 9.1 Classification Metrics
In multi-class HGR problems—especially when dealing with class imbalance—researchers often report both micro-averaged (per-instance) and macro-averaged (per-class) versions of common classification metrics. Micro-averaged metrics aggregate contributions from all instances across classes, thus being sensitive to class imbalance. Macro-averaged metrics compute the metric independently for each class and then average, treating all classes equally regardless of their frequency. The formulae for each metric can be found in the Table 10.
Table 10: Common utilized metrics for evaluation of hand gesture classification. Here, $TP_{i}$, $FP_{i}$, $TN_{i}$, and $FN_{i}$ denote the number of true positives, false positives, true negatives, and false negatives for the $i$-th class, respectively, and $N$ is the total number of classes.
Metrics Micro-averaged formulae Macro-averaged formulae
Accuracy $\frac{\sum_{i=1}^{N}(TP_{i}+TN_{i})}{\sum_{i=1}^{N}(TP_{i}+TN_{i}+FP_{i}+FN_{i})}$ $\frac{1}{N}\sum_{i=1}^{N}\frac{TP_{i}+TN_{i}}{TP_{i}+FP_{i}+TN_{i}+FN_{i}}$
Precision $\frac{\sum_{i=1}^{N}TP_{i}}{\sum_{i=1}^{N}(TP_{i}+FP_{i})}$ $\frac{1}{N}\sum_{i=1}^{N}\frac{TP_{i}}{TP_{i}+FP_{i}}$
Recall $\frac{\sum_{i=1}^{N}TP_{i}}{\sum_{i=1}^{N}(TP_{i}+FN_{i})}$ $\frac{1}{N}\sum_{i=1}^{N}\frac{TP_{i}}{TP_{i}+FN_{i}}$
Jaccard-Index $\frac{\sum_{i=1}^{N}TP_{i}}{\sum_{i=1}^{N}(TP_{i}+FP_{i}+FN_{i})}$ $\frac{1}{N}\sum_{i=1}^{N}\frac{TP_{i}}{TP_{i}+FP_{i}+FN_{i}}$
F1-Score $2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$
@@ 9.2 Sequence-to-Sequence Metrics
For continuous HGR (CHGR), particularly in Continuous Sign Language Recognition (CSLR), sequence-level metrics are used. Word Error Rate (WER) (VAC-CSLR; ADADOGLOU) is the standard metric, measuring the minimum number of substitutions ($S$), deletions ($D$), and insertions ($I$) needed to transform the predicted sequence into the ground truth sequence of length $N$:
@@
Edit Distance Accuracy (FA-STGCN; CTCX_WANG2022123) is a complementary metric representing the ratio of correctly predicted elements in the sequence:
$\text{Edit-Accuracy}=\frac{N-S-D-I}{N}$ (2)
@@ 9.3 Computational Resource Metrics
  • 3.
    Model Size in Megabytes (MBs): The total memory footprint of the model.
  • 4.
    Number of Parameters: The number of trainable parameters.
  • 5.
    Floating Point Operations (FLOPs): Measures the number of numerical operations required to run a model (either during training or inference time). Being independent of the hardware utilized, it could be considered - along with the number of parameters - more reliable.
@@ 10 Current Challenges and Future Research Directions
As mentioned in the Methods sections (Sections 4, 5, and 6), researchers have developed sophisticated techniques and large-scale datasets to achieve satisfactory performance. However, as shown in the relevant tables (Tables 2, 3, 4), recognition accuracy remains low in many cases—particularly in more complex tasks such as CSLR. In this section, we discuss the current challenges and limitations that hinder efficient VHGR, posing ongoing concerns for the research community and shaping future research directions. Additionally, we review the techniques proposed to address these issues, offering insights that may guide future efforts to overcome the identified challenges.
@@ 10.1 Gesture-Irrelevant Factors
The presence of gesture-irrelevant factors (or redundant information) hinders accurate recognition in diverse real-world scenarios. Some of these factors include the following (EDenseNet; Mohammad_Mahmudul_static_hgr; 2DPSTPP; ULTRA-RANGE-UGV-Bamani; EnGesto; RAAR3DNet):
  • 1.
@@
Aiming to address these challenges, researchers have proposed segmentation models to crop hand regions and remove background interference (FGDSNet; tuan_linh_pham_dang2023lightweight; DRCAM_TIM), object detectors for spatial normalization (ULTRA-RANGE-UGV-Bamani; Mohammad_Mahmudul_static_hgr), super-resolution techniques to enhance image quality (ULTRA-RANGE-UGV-Bamani), attention mechanisms (RAAR3DNet; Benjia_Zhou) to enforce the visual encoder to pay attention on the most informative regions, information-theoretic losses (Yunan_Li), and feature selection techniques (SeST) to reduce information redundancy. Many of these approaches rely on computationally expensive off-the-shelf models, which increase the complexity of the overall HGR methods and set an upper bound to their accuracy.
The approaches mentioned previously may be enhanced by using eXplainable Artificial Intelligence (XAI) (Rodis2024Access) techniques, which could be adopted to understand better the behavior of the networks. For instance, Grad-CAM could be further utilized to determine the image region that affects more the predicted gesture class (as has already been done by a few researchers: (CorrNet; SWIN_MSTP_Neurocomputing_2025_CSLR)). Such results may indicate the effectiveness of the embedded on the network attention mechanisms or other related modules to emphasize the most important parts of the input image. t-SNE visualizations could also be leveraged to evaluate the quality of the extracted features, and in particular, how much they are influenced by the presence of redundant information. Thus, guided by the tools provided by the XAI or other similar framework, the interference of gesture-irrelevant factors could be mitigated through the careful design of the future architectures.
10.2 Computational Cost
Even lightweight methods for SHGR and IDHGR achieve high accuracy. Although, for the more complex task of CHGR, the cost is essential; all RGB-based methods (that are significantly more accurate than pose-based) require more than 366.3 GFLOPs ((DSTEN-CSLR-yu2025dsten)) and reach up to 2950.5 GFLOPs ((TemporalLiftPooling)). This issue prevents not only the deployment of such methods but also the efficient training, which takes longer. Pose-based methods (CoSign) are significantly lighter, but the pose estimation cost is not taken into account. The number of parameters does not change dramatically between IDHGR and CHGR methods, while many SHGR methods require more parameters (e.g. Gesture-CNN (G-CNN-SHARMA2021115657) has 67.25M parameters).
The need of large computational resources does not only prevent the real-time and commercial deployment of VHGR systems, but also put barriers to the evolution of the field, as the conducted experiments take many hours or even days to complete. To mitigate this, authors have focused on reducing the complexity of their models by using non-parametric layers (GestFormer) or designing smaller architectures (LHGR-Net). Thus, reducing the total complexity of the models, could be further examined in the future. Considering also recently proposed approaches for DL models compression (DL_MODEL_COMPRESSION_marino2023deep), such as pruning (i.e., removing less important weights) and quantization (i.e., reducing the arithmetic precision of the parameters), may contribute to this direction. Freezing layers of pretrained models could also be investigated in order to reduce the time required for fine-tuning.
@@ 10.3 Data Scarcity
Many existing datasets either contain a relatively small number of samples or lack sufficient diversity. For example, among static HGR datasets, only the versions of HaGRID (WACV_kapitanov2024hagrid; HAGRIDv2_ARXIV) have been collected from more than one hundred participants, despite some others having a plausible number of samples (LaRED; ULTRA-RANGE-UGV-Bamani). A similar situation is observed in IDHGR datasets, where only the Jester dataset (Jester-Dataset) includes data from over one hundred subjects. Furthermore, the amount of samples for specific modalities, such as IR data, is limited compared with the size of RGB-based datasets, an issue known as data imbalance (IR_data_imbalance_mobilenet). The lack of large-scale, diverse datasets leads to overfitting and limited generalization, which is closely related to the challenges discussed earlier. This issue is further exacerbated in tasks that involve costly annotation processes and require data-intensive deep learning models with a large number of parameters, such as CSLR.
@@
  • 1.
    Multilingual Data Aggregation: Combining datasets (e.g., CSL and DGS) to exploit cross-lingual visual similarities increases the overall amount of training data (Fangyun_Wei_ICCV).
  • 2.
    Self-Supervised Learning: In the context of the self-supervised learning setting (konstantakos2025self), gloss-to-gloss translation tasks have been used for parameter initialization (CVT-SLR), providing linguistic priors to sign language models. Pretraining with pose reconstruction has also been investigated to incorporate prior knowledge of hand structure (SignBERT+; SignBERT_ICCV; omar_ikne).
  • 3.
    Advanced Data Augmentation: Adversarial learning has been employed to optimize augmentation hyperparameters (ADVERSARIAL_DATA_AUGM), making the models robust to spatial deformations.
  • 4.
    Few-/Zero-shot Learning: Approaches using only a small number of samples to correctly recognize gestures—often based on CLIP—have also been proposed (ZSSLR; FSSLR; ZS-DHGR; ZS-GR; ZERO-SHOT-SLR).
Generative AI, particularly diffusion models (DMs) for image augmentation (alimisis2025advances), could help address data scarcity in HGR. Existing datasets are often limited in size or focused on common tasks, like well-known sign languages. DMs can increase dataset diversity by generating additional samples with varied scenes, subjects, and lighting, or support specialized domains such as rare sign languages and human-robot interaction in firefighting or military scenarios. Thus, leveraging DMs represents a promising direction for future work.
@@ 10.4 Optimization Difficulties
Training the visual encoder effectively remains difficult, especially in deep architectures for CSLR. As shown in CTCA (CTCA), (Fig. 9), increasing the depth of the temporal module (e.g., 1D-TCN followed by BiLSTM) improves its generalization but hampers the encoder’s learning due to gradient attenuation in backpropagation. Furthermore, the CTC spike phenomenon—where the model over-focuses on a few keyframes—can cause rapid overfitting and limit encoder training (VAC-CSLR; SMKD).
Figure 9: Experimental results in (CTCA) analyze the effect of the temporal aggregation module’s depth—a 1D-TCN with varying layers followed by a two-layer BiLSTM—on training the visual encoder (spatial perception module). In both charts, the x-axis shows the number of 1D-TCN layers. The top chart plots information stored in weights (IIW) (pac-bayes-wang2022pacbayes), indicating generalization ability, while the bottom shows WER. Increasing 1D-TCN layers improves short-term temporal generalization but reduces overall performance due to undertrained visual encoder weights.
@@
  • 1.
    Auxiliary Supervision: Adding early-stage classifiers to the encoder or short-term module has been proposed as a strategy to guide training by providing additional supervision (VAC-CSLR; CSGC_Rao_Sun_Wang_Wang_Zhang_2024; CorrNet; corrnetplus_hu2024corrnetsignlanguagerecognition; TCNet_Lu_Salah_Poppe_2024; C2SLR).
  • 2.
    Temporal Module Redesign: Placing the 1D-TCN and BiLSTM in parallel instead of sequentially (CTCA).
  • 3.
    Linguistic Constraints and Gloss Segmentation: Using gloss boundaries (SMKD) or prior knowledge (GPGN) to regularize learning.
The aforementioned approaches —particularly auxiliary supervision (VAC-CSLR)— have served as the foundation for many existing methods. However, the increasing complexity of modern deep architectures often exceeds the effectiveness of these techniques. As a result, optimization remains a significant challenge —one that may be further exacerbated by the continued development of even deeper models.
@@ 10.5 Limitations of Architectures
DL methods for HGR have increased the recognition performance. However, the existing architectures suffer from limitations, including the following:
  • 1.
    Block-fixed models: The existing architectures are based on hand-crafted design, consisting of block of layers with hyperparameters (e.g., number of channels, convolution kernel size) that are defined manually by the researchers. This approach is suboptimal (RAAR3DNet), as several other possible configurations have not been explored.
  • 2.
    Insufficient graph representations of GCNs: The standard GCNs utilized fixed topology graphs, often based on the shape of the hand, to better capture the interactions between the various skeleton joints. However, this configuration has been proved to be suboptimal (TD-GCN), unable to efficiently capture the temporal dynamics.
  • 3.
    Ineffective multimodal fusion: Out of the existing multimodal methods, most of them perform late fusion to incorporate the information of multiple streams (SignBERT_ICCV; SignBERT+; BEST; Weichao-Zhao; MASA; SKIM; SAM_SLR; jiang2021sign-sam-slr-v2). This scheme does not fully leverage the available modalities, as the features extracted by each stream are isolated from the others during the processing on the backbones. Thus, the obtained vector representations do not benefit from the utilization of multiple modalities.
To address these limitations, researchers have proposed various strategies, including Neural Architecture Search (NAS) to identify optimal model structures through searching on a large set of configurations (RAAR3DNet), learnable or time-dependent skeleton adjacency matrices that determine implicitly the optimal way to integrate the features of each joint (DSTA_SLR_hu2024dynamic; DSTSA-CGN_cui2025dstsa; ST-PSM_L-PSM; TD-GCN) and multi-level fusion for dual-stream approaches, which enhances the quality of the extracted features by early incorporating information from various sources (CTFB_Hampiholi).
Despite these advances, several aspects of deep learning architectures remain underexplored, leaving a substantial research gap. For instance, to the best of our knowledge, multi-level and mid-level multimodality fusion have been investigated only in the context of dual-stream approaches. Thus, future work could consider the intermediate fusion of three or more streams, e.g. using fusion modules (MMTM; CTFB_Hampiholi) and identifying the optimal way to connect the different streams.
@@ 11 Conclusions
The undeniable importance of Vision-based Hand Gesture Recognition (VHGR) stems from its wide range of applications—spanning human-robot interaction to sign language understanding. This broad applicability has made it a vibrant and active research field. Recent advances in deep learning (DL) have played a pivotal role in driving progress, enabling the development of increasingly sophisticated architectures and training strategies aimed at tackling more complex and fine-grained tasks, such as continuous sign language recognition (CSLR).
This survey focused on VHGR methods that leverage visual input and DL. These methods were highlighted not only for their superior performance compared to traditional hand-crafted approaches, but also for their higher potential for deployment in real-world scenarios, unlike more cumbersome sensor-based techniques. From the surveyed literature, it becomes clear that even relatively simple and shallow DL models can achieve high accuracy in static HGR tasks. In such cases, the primary challenges are typical of general computer vision problems and are not inherently tied to HGR itself. For isolated dynamic and continuous VHGR more advanced approaches have been proposed, aiming to capture more complicated context. The complexity of the architectures (especially for CSLR) cause serious optimization difficulties, leading researchers to develop sophisticated training setups.
Nevertheless, the most important challenges regarding VHGR are not yet solved, leaving substantial research gaps; the data scarcity - especially concerning more specialized tasks, such as the recognition of rare sign languages - and the high computational cost remain essentially unresolved. Furthermore, the existing methods are limited to a predefined list of gestures, which may not be optimal for real-world interaction with humanoid robots or interpretation of the body language. Thus, VHGR could be expected to open-vocabulary settings, enabling deployment to broader applications. The potential of leveraging the extensive knowledge of foundation models (kas2025foundation_models) in the context of VHGR enabling zero-shot gesture recognition as well as the utilization of model compression techniques (DL_MODEL_COMPRESSION_marino2023deep) to reduce the computational complexity are also rarely explored, leaving space for future work.
Acknowledgments
The research leading to these results received funding from the European Commission under Grant Agreement No. 101168042 (TRIFFID) and No. 101189557 (TORNADO).